Chris walks 3/8 mile each day to school.
Anna walks 1/2 mile . Do they walk the same distance to school? Explain

Answer by TimothyLamb(4379)

---
3/8 ?=? 1/2
---
find a common denominator
in this case 8 can be a common denominator for both fractions:
---
3/8 ?=? (1/2)(4/4)
3/8 ?=? (1*4)/(2*4)
3/8 ?=? (4)/(8)
3/8 ?=? 4/8
---
3/8 does not equal 4/8
---
answer:
Chris and Anna do NOT walk the same distance to school
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php